How can I solve this distributive problem???
(104)(9)

Answer by stanbon(75887)

How can I solve this distributive problem???
(104)(9)
-------------
I think what they want you to do is show how you could
use the distributive law to evaluate the product.
--------------
For example:
(104)(9) = (100 + 4)9 = 900 + 36 = 936
============================
Cheers
Stan H.